Circular No. 8283

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004S IN MCG –05-16-21 N. Suntzeff, Cerro Tololo Interamerican Observatory (CTIO), reports

that spectrophotometry (range 310–950 nm; resolution 400) of SN 2004S (cf. IAUC 8282) was obtained by A. Globus, L. Galli, A. Whiting, and L. Schmidtobreick at the CTIO 1.5-m telescope (+ facility spectrograph) on Feb. 6.1 UT. The spectrum of the supernova is that of a type-Ia event, around 5 days after maximum light. The expansion velocity (relative to the galaxy redshift of 2731 km/s) of the Si II 635.5-nm absorption line is ∼ 9300 km/s.

K. Krisciunas, Department of Physics, University of Notre Dame, in collaboration with the SMARTS consortium, reports photometry of SN 2004S obtained by J. Espinoza at the CTIO 1.3-m reflector on Feb. 7.225 UT (calibration via the Landolt field Rubin 149; internal errors ± 0.01 mag): U = 14.64, B = 14.55, V = 14.45, R = 14.39, I = 14.76.

SUPERNOVA 2004Q IN ESO 507-G11 M. Hamuy, Las Campanas Observatory, reports that a spectrum (range

360–900 nm) of SN 2004Q (cf. IAUC 8281), obtained on Feb. 5.30 UT with the Clay 6.5-m telescope (+ LDSS-2), shows that this is a young type-II supernova characterized by a blue continuum with low-contrast Hα, Hβ, and He I 587.6-nm lines exhibiting P-Cyg profiles.

COMETS C/2003 E2, 2003 E3, 2003 E4, 2003 F3, 2003 F4 (SOHO) Further to IAUC 8278, K. Battams reports measures of additional

Kreutz sungrazing comets found on SOHO C3 website images by D. Evans (C/2003 E2), S. Hönig (C/2003 E3), J. Sachs (C/2003 E4), M. Oates (C/2003 F3), and X. Leprette (C/2003 F4). C/2003 F3 and C/2003 F4 were also detected in C2 images. The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2004-C31.

Comet  2003 UT  α_{2000}  δ_{2000}
C/2003 E2 Mar. 4.154  23^{h}14.^{m}3  − 7^{o}28′
C/2003 E3  11.446  23 41.1  − 4 28
C/2003 E4  15.529  23 55.5  − 2 51
C/2003 F3  20.196  0 20.6  − 1 26
C/2003 F4  25.404  0 29.7  + 1 03
2004 February 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green